OpenStudy (anonymous):

what is the order of 1.5x10 ^ 3,1.4x10-1,2x10 ^3, and 1.4 x 10-2 from least to greatest?

OpenStudy (anonymous):

and 1.4 x 10-2; 1.4x10-1;1.5x10 ^ 3,,2x10 ^3,

OpenStudy (unklerhaukus):

\[ 1.4 \times 10^{-2}=0.014\]\[1.4\times10^{-1}=0.14\]\[1.5\times10^ 3=1500.\]\[2\times10 ^3=2000.\]